How much heat energy does it take to boil 2.35 kg of water at 100oC completely to steam at 100oC ?
Answer by ikleyn(52834)

.
How much heat energy does it take to boil 2.35 kg of water at 100oC completely to steam at 100oC
~~~~~~~~~~~~~~~~~~


Here the temperature is the same of 100°C.



Only latent heat of vaporization of steam is required to convert water to steam

    Q = m * L

where  m is the mass of water  = 2.35 kg;

       L is the latent heat of vaporization = 2257 kJ/kg



The required heat energy = 2.35 * 2267 = 5327.45 kJ.      ANSWER


Solved.